If Ben invests $4000 at 4% interest per year, how much additional money must he invest at 5 1/2% annual interest to ensure that the interest he recieves each year is 4 1/2% of the total amount invested?
Answer by stanbon(48551)

If Ben invests $4000 at 4% interest per year, how much additional money must he invest at 5 1/2% annual interest to ensure that the interest he recieves each year is 4 1/2% of the total amount invested?
------------------
Let "x" be the amount of additional money.
EQUATION:
interest + interest = interest
0.04(4000)+0.055x = 0.045(4000+x)
160+0.055x=180+0.045x
0.01x=20
x=$2000 (amount of additional money he must invest)
Cheers,
Stan H.